OpenStudy (anonymous):

Which slope lies on the line with point slope equation y+5= 2(x+8)? A. (8,-5) B. (-8,5) C. (-8,-5) D. (8,5)

OpenStudy (texaschic101):

y + 5 = 2(x + 8) In point slope form, y - y1 = m(x - x1)......x1 = 8 and y1 = 5 (8,5) ....however, this is not a slope...these are just points that lie on the line